In ANOVA, if the F Ratio = 1.85 and the F critical boundary (from the tables) = 3.55, what conclusion should be drawn?

all the means are different

all the means are equal

at least one of the means is different

all the standard deviations are different

What is the correct conclusion based on the following ANOVA result: [F(3,16) = +2.24, p > .05]

dfbn = 16; dfwn = 3; k = 17; result was NOT significant

dfbn = 3; dfwn = 16; n = 4; result was NOT significant

dfbn = 3; dfwn = 16; k = 4; result was NOT significant

dfbn = 16; dfwn = 3; n = 17; result was NOT significant
Answer #1

Solution:

1)

F ratio = 1.85 <  F critical 3.55

So , fail to reject H0 : all means are equal.

So , conclusion is

all the means are equal

2)

F(3,16) = +2.24

dfbn = 3

dfwn = 16

But dfbn = k - 1

k = dfbn + 1 = 3 + 1 = 4

So , correct option is

dfbn = 3; dfwn = 16; k = 4; result was NOT significant